OpenStudy (anonymous):

How do you solve: sqrt(16x^5)

OpenStudy (anonymous):

\[\sqrt{16x^5}\]

sam (.sam.):

You mean simplify right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

4x^(5/2)

sam (.sam.):

\[\sqrt{16x^5} \\ \\ \text{Since 16 and x is multiplying} \\ \\ \sqrt{16}\sqrt{x^5} \\ \\ 4\sqrt{x^5}\]

OpenStudy (anonymous):

Sam, I was taught that you could simplify the X too down to 4x^2\[\sqrt{x}\]

OpenStudy (anonymous):

So what do you do with the x?

sam (.sam.):

No you can't simplify that, because \[4\sqrt{x^5}=4x^{\frac{5}{2}}\]

sam (.sam.):

square root is half of an exponent

OpenStudy (anonymous):

OOOOOOOOHHHHH ok thank you soooooo much. :D

sam (.sam.):

@KowaiPanda welcome :)